# Data-Driven Room Acoustic Modeling Via Differentiable Feedback Delay Networks With Learnable Delay Lines

### v2

@@ 5.3 Parameter Initialization
Figure 6: Gym (h052) IRs.
5.4 Implementation Details
We implement our differentiable model in Python using PyTorch. We define the FDN as a class inheriting from nn.Module. We thus define the unconstrained trainable parameters as instances of nn.Parameter. Our model operates at a sampling rate of 16 kHz. As a result, its memory footprint turns out to be contained, allowing us to train all FDNs considered in the present study on a single NVIDIA Titan V graphics card. [^5] Preliminary experiments carried out with increased computational resources indicate that results comparable with those reported in the present study can be obtained at a sampling rate of 48 kHz. We optimize the models for a maximum of 650 iterations using Adam [42] with a learning rate of $0.1$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and no weight decay. The average training time per iteration is reported in Table 1. At each step, approximately 25% of the time is taken by the forward pass of the FDN, just below 20% is spent computing the loss function, and just above 55% is spent backpropagating the gradients and updating the parameters. Notably, we observe that the computation time increases linearly with the estimated $T_{60}$. For each test case, we present the model with the lowest composite loss.
5.5 Test Case: Gym (h052)
We start by considering the Gym RIR (h052). The FDN is trained setting $\lambda=1$, i.e., equally weighting $\mathcal{L}_{\text{EDC}}$ and $\mathcal{L}_{\text{EDP}}$. As shown in Table 1, the best model is reached at iteration 279, after the loss has decreased by two orders of magnitude with respect to the initial value obtained with the random initialization described in Section 5.3. Here, we present a comparison between the target room acoustics (solid black line), the GA baseline [16] (dotted blue line), the HRTC baseline (dash-dotted green line), the Colorless baseline [18] (dash-dotted blue line), and, finally, the proposed differentiable FDN (dashed orange line).
Figure 5, 5, and 6 show that the proposed method is capable of closely matching the EDC, EDP, and envelope of the target RIR, respectively.
@@ 5.5 Test Case: Gym (h052)
Figure 9: Hallway (h270) IRs.
Conversely, the baseline methods produce poorer results. In Figure 5, we may notice, in fact, that the EDC of the GA baseline deviates from that of the target RIR after just below 150 ms. The energy decay of the HRTC method, instead, is matched just for the very first few ms, showing to part from the target immediately after and then linearly decaying. It is interesting to notice that the Colorless method, although sharing some parameters with HRTC, behaves completely different. In fact, after an initial overshoot, said baseline is able to match the target decay till 400 ms, where it starts to exhibit a less steep decay. Such a difference is due to the feedback matrix and input/output gains being optimized to get a colorless IR, unveiling their importance for obtaining a correct time decay.
Furthermore, the EDP of the baseline methods indicate a scarce echo density in the first 250 ms, especially for HRTC and Colorless. The output of the FDN obtained with GA, instead, becomes identically zero just after approximately 0.9 s. The ensuing EDP pathologies are confirmed by the IR depicted in Figure 6, where the baselines are shown to yield fewer, more prominent taps compared to the IR of the proposed FDN model.
Notably, however, Figure 6 also suggests that both the proposed and baseline methods estimate a larger $d$ than what would correctly render the direct sound. We attribute this phenomenon to an attempt at compensating the lack of a noise floor that, in real-life measurements, contributes to the total energy of the RIR. We argue that, in FDN models with tunable direct gain, offsetting this bias is naively achieved by increasing $d$.
As far as reverberation metrics are concerned, Table 4 shows that the proposed method has an overall better performance, with a $\Delta T_{30}$, $\Delta D_{50}$ and $\Delta t_{s}$ of 2.5 ms, 0.73%, and 59 µs, respectively. The $T_{20}$ is estimated with an error of 100 ms, i.e., approximately 2.3 and 4 times lower than those of the GA and HRTC methods, respectively. As for this metric, our method is second only to the Colorless method, whose $\Delta T_{20}=15$ ms. Similarly, as far as $T_{60}$ is concerned, the proposed approach is surpassed only by the HRTC method that shows a $\Delta T_{60}=6.8$ ms; this is expected given that the parameter $\gamma$ is specifically designed to match the target $T_{60}$. Finally, our method shows the second best $\Delta C_{80}$ overcome only by GA’s, with a difference of just 0.13 dB.
@@ 5.6 Test Case: Hallway (h270)
Figure 12: Pizzeria (h214) IRs.
Let us now consider the Hallway RIR (h270), which is characterized by nearly half the $T_{60}$ of the previous case. Table 1 shows that the best results are obtained at iteration 340, where the loss is again lower by two orders of magnitude with respect to the starting point.
Figure 8 reports the EDCs of target, baselines, and proposed methods following the color conventions reported in the previous subsection. Once again, we evince the good matching between our and target decays, as they start to part around $0.4$ s where the residual energy of the predicted RIR is low, i.e., around $-40$ dB. Instead, HRTC and Colorless methods present a sharp drop of energy after the direct path, which results into IRs characterized by an almost total absence of reflections for the first few ms, as shown in Fig. 9. Then, HRTC presents an overshoot that lasts till 3̃50 ms, after which it starts to match the target EDC. However, this happens when the energy has already reached $-40$ dB, thus after virtually all early reflections have occurred. In addition, the Colorless and GA baselines are not able to correctly match the decay in time except for a small span of time around $100$ ms.
The EDPs depicted in Figure 8 indicate that our approach is able to generate a close match with the target EDP, further validating the effectiveness of proposed Soft EDP loss function. Indeed, the orange dashed curve nicely follows the target curve until the $T_{60}$; we remind, in fact, that the training is performed only on such a span of time. Instead, the EDP of the baseline methods show low values in the first 200 ms compared to the target one, and afterward, even if they take on comparable values, they do not follow the same trend. This is confirmed by the RIRs shown in Figure 9. Indeed, even in this test case, the FDN optimized by means of the proposed approach has an IR that resembles more of a realistic RIR, even though the amplitude of individual taps is not entirely matched.
Table 4 reports the reverberation metrics. We can observe that the proposed optimization approach yields better results with respect to the previous case. Notably, the proposed optimization performs better than the three baselines for all the metrics but the $T_{60}$, for which, as expected, the HRTC method shows the lowest error, with a $\Delta T_{60}$ of $11$ ms against the $95.6$ ms obtained with our approach. Then, the error $\Delta C_{80}$ associated with the clarity index is three orders of magnitude less than what can be achieved with the GA and HRTC methods, and two orders of magnitude less than Colorless’. In addition, the error $\Delta t_{s}$ on the center time reaches 32 µs, while all the baselines have a greater error, shifting the center of mass of the predicted IR energy more toward the reverberation tail.
@@ 5.7 Test Case: Pizzeria (h214)
Finally, let us address the shortest RIR of the ones considered in the present study, i.e., h214, having a $T_{60}$ of just above 0.2 s.
Here, all the baseline methods appear to fail at modeling the target room. In particular, the IRs shown in Figure 12 exhibit a few sparse taps, being thus far from resembling a real RIR. Since the backward-integrated energy abruptly decays with every peak, this results in a staircase-like behavior of the EDC depicted in Figure 11.
Facing the same difficulties, the proposed optimization method takes more gradient steps compared to the previous two test cases before converging at iteration 583.
In Figure 11, the resulting EDC (dashed orange line) closely follows that of the target RIR (solid black line) until the two reach approximately $-47$ dB. Then, the two deviate before meeting again at $-60$ dB. Still, a direct comparison between EDCs in the range of $-50$ dB to $-70$ dB is not entirely reliable since the RIR has so little energy that background and sensor noise take on a much more relevant role when it comes to integrating the energy of the measured signal.
Overall, the proposed optimization method proves to perform well in fitting the RIR under scrutiny, with Table 4 reporting an error on all three reverberation time metrics of approximately 10 ms, $\Delta D_{50}$ of 0.016%, and $\Delta t_{s}$ of 19 µs. As in the previous test case, the only metric for which the proposed method shows a lower performance with respect to one of the baselines is $T_{60}$. As a matter of fact, the HRTC is the only baseline to be once again characterized by a lower $\Delta T_{60}$, with a value of $6.9$ ms.
@@ 5.8 Excluding the EDP Loss Term
In developing our method, we noticed that using only the EDC loss function leads to ill-behaved IRs. Namely, we found that, while closely matching the desired EDC, the IR of an FDN trained using only $\mathcal{L}_{\text{EDC}}$, i.e., $\lambda=0$, tends to exhibit an unrealistic echo distribution compared to the RIRs of real-life environments. In this section, we compare the results of our differentiable FDN trained without EDP regularization loss with those presented in Section 5.6 obtained using the composite loss function in (6). For conciseness, we limit our analysis to the Hallway RIR (h270); results obtained with other RIRs are comparable to what is shown below.
When excluding the EDP loss term from (6), the NMSE between true and predicted EDCs (8) is slightly lower than that of the proposed method, i.e., $4.7\times 10^{-3}$ and $5.2\times 10^{-3}$, respectively. Yet, the MSE between true and predicted EDPs (15) is $0.395$, i.e., two orders of magnitude grater than the $0.0046$ reported in Table 1 for the proposed method. This can be observed in Figure 13b, showing that the echo density when $\lambda=0$ (dash-dotted purple line) is far from the desired profile (solid black line). The target RIR and the proposed method (dashed orange line) produce an EDP with values consistently around one, indicating dense reverberation. On the contrary, the FDN trained without Soft EDP regularization yields an EDP with small values, signaling an uneven echo density due to the presence of a few prominent reflections [38]. This observation is confirmed by the IR shown in Figure 13a that exhibits a few clearly noticeable taps peaking above a denser reverberation tail.
Excluding the EDP loss term means solving a minimization problem with no constraints discouraging the model to use just a small number of delay lines to capture the overall EDC behavior. Since FDNs extend a parallel comb-filter structure, we believe that the behavior observed in Figure 13a is related to the well-understood problem occurring when short delay lines are strongly excited, which, in turn, aggravates the comb-like behavior of the FDN, ultimately resulting in an unpleasing metallic sound quality [41, 18].
@@ 5.9 Soft EDP Approximation
Finally, we discuss the approximation capabilities of the proposed Soft EDP function introduced in Section 4.5.2. Figure 14 shows the non-differentiable EDP (solid black line) defined in (9) against several Soft EDP approximations of the three RIRs considered in the present study. We test various scaling parameters $\kappa$, namely $10^{2}$, $10^{3}$, and $10^{4}$, along with the proposed time-varying $\kappa_{n}$ linearly increasing from $10^{2}$ ($n=0$) to $10^{5}$ ($n=L_{T_{60}}-1$). We depict the profiles only for time indices below the $T_{60}$, as this range is the one considered when training the FDNs.
@@
HRTC $1.2592$ $0.3977$ $1.2526$ $0.3365$ $1.2189$ $\underline{0.0068}$
Colorless $0.8769$ $\underline{0.0154}$ $1.3036$ $0.3875$ $1.3551$ $0.1294$
Ours $0.7607$ $0.1009$ $0.9136$ $\underline{0.0025}$ $1.1854$ $0.0403$
@@
Target $12.1061$ $89.0091$ $22.8994$
GA $11.9853$ $\underline{0.1208}$ $91.1939$ $2.1849$ $19.3434$ $3.5560$
HRTC $7.8383$ $4.2678$ $79.4541$ $9.5550$ $38.1226$ $15.2232$
Colorless $10.1037$ $2.0024$ $80.4545$ $8.5546$ $27.8442$ $4.9449$
Ours $11.8266$ $0.2795$ $88.2766$ $\underline{0.7325}$ $22.9584$ $\underline{0.0591}$
Table 3: Metrics for the Hallway RIR ($\lambda=0.1$).
@@
HRTC $0.6341$ $0.1052$ $0.6325$ $0.0315$ $0.6178$ $\underline{0.0111}$
Colorless $0.4350$ $0.0939$ $0.4596$ $0.1414$ $0.5730$ $0.0337$
Ours $0.5031$ $\underline{0.0259}$ $0.5716$ $\underline{0.0294}$ $0.7023$ $0.0956$
@@
Target $14.0797$ $90.6917$ $20.1163$
GA $15.5019$ $1.4223$ $92.1064$ $1.4147$ $18.3649$ $1.7514$
HRTC $11.4092$ $2.6705$ $88.2675$ $2.4242$ $20.8044$ $0.6881$
Colorless $13.7608$ $0.3188$ $87.4117$ $3.2801$ $18.7276$ $1.3887$
Ours $14.0713$ $\underline{0.0083}$ $90.5123$ $\underline{0.1794}$ $20.1487$ $\underline{0.0324}$
Table 4: Metrics for the Pizzeria RIR ($\lambda=0.1$).
@@
HRTC $0.2252$ $0.0609$ $0.2151$ $0.0358$ $0.2131$ $\underline{0.0069}$
Colorless $0.2288$ $0.0645$ $0.2050$ $0.0256$ $0.1911$ $0.0151$
Ours $0.1543$ $\underline{0.0099}$ $0.1694$ $\underline{0.0100}$ $0.2166$ $0.0104$
@@
Target $30.9883$ $99.3819$ $7.1698$
GA $40.6833$ $9.6950$ $99.7989$ $0.4170$ $6.4524$ $0.7174$
HRTC $25.1428$ $5.8455$ $98.6346$ $0.7473$ $9.4239$ $2.2541$
Colorless $23.9120$ $7.0763$ $97.8538$ $1.5281$ $10.0684$ $2.8986$
Ours $32.1683$ $\underline{1.1801}$ $99.3988$ $\underline{0.0168}$ $7.1891$ $\underline{0.0194}$
Figure 13: (a) IR and (b) EDP obtained by training the FDN without EDP regularization loss ($\lambda=0$).
@@ 5.10 Limitations
In this work, we focused on two important perceptual characteristics of room impulse responses: energy decay and echo density. While rendering these acoustic features is essential for any artificial reverberation algorithm that aims to be realistic, by themselves, energy decay and echo density are not enough to comprehensively model all perceptual qualities of reverberation. In particular, matching the spectral envelope of the measured RIR turns out to play a crucial role in room acoustic simulation [39].
However, Figures 17, 17, 17 show that none of the FDNs considered in the present study manages to capture the frequency response of the target RIRs. This, after all, is not among the objectives of the respective parameter selection/tuning algorithms, except for the GA baseline, whose fitness function amounts to a $L^{1}$-loss between MFCCs. Still, despite considering cepstral features, the magnitude of the resulting transfer function is far from resembling the target spectrum.
On the one hand, such a conspicuous mismatch entails that the output signals of the FDNs sound different not only from one another but also from the corresponding input signal convolved with the target RIR. On the other hand, to the best of our knowledge, there is no easy way to decouple spectral coloration from sound energy decay without affecting the latter.
In turn, this undermines the reliability of any perceptual test assigning similarity scores to each method with respect to the target, as subjective judgments would be significantly influenced by differing coloration. In this respect, pilot experiments proved inconclusive, highlighting the need for further investigation into spectro-temporal modeling.
In practice, it would be possible to extend the cost function to the frequency-dependent case by including an additional regression loss term in (6). However, a thorough analysis of this approach lies beyond the scope of the present work and is left for future work.

### v3

@@ 5.5 Test Case: Gym (h052)
Figure 6: Gym (h052) IRs. The time axis is limited to the $T_{60}$ for visual clarity.
@@
Figure 9: Hallway (h270) IRs. The time axis is limited to the $T_{60}$ for visual clarity.
@@ 5.6 Test Case: Hallway (h270)
Let us now consider the Hallway RIR (h270), which is characterized by nearly half the $T_{60}$ of the previous case. Table 1 shows that the best results are obtained at iteration 796, where the loss is again lower by three orders of magnitude with respect to the starting point.
Figure 7 reports the EDCs of target, baselines, and proposed methods following the color conventions reported in the previous subsection. Once again, we evince the good matching between ours and the target decay, especially in the early and late portion of the curve. GA correctly matches the target EDC only in the first $100$ ms before rapidly decaying. The HRTC and Colorless methods, instead, present a sharp energy drop after the direct path, which results in IRs characterized by an almost total absence of reflections for the first few ms, as shown in Figure 9. HRTC, in turn, presents a slight overshoot that lasts for the first 300 ms, after which it closely matches the target EDC. The proposed method, instead, deviates from the target in its central part, approximately from 180 to 500 ms. Arguably, however, these two kind of errors are not equivalent since the earlier portion of the RIR is known to be more relevant from a perceptual point of view [38]
Figure 8 indicates that our approach is able to closely match the target EDP, further validating the effectiveness of the proposed Soft EDP loss function. Indeed, the orange dashed curve follows the target curve until the $T_{60}$. We remind, in fact, that the training is performed only on such a span of time. Conversely, the EDP of the baseline methods show low values in the first 200 ms compared to the target one, and afterward, even if they take on comparable values, they do not follow the same trend. This is confirmed by the RIRs shown in Figure 9. Indeed, even in this test case, the FDN optimized by means of the proposed approach has an IR that resembles more of a realistic RIR, even though the amplitude of individual taps is not entirely matched.
Table 3 reports the reverberation metrics. We can observe that the proposed optimization approach yields results comparable to the previous case, outperforming the baseline methods in five out of the six metrics. This time, due to the aforementioned mismatch in the central part of the EDCs in Figure 7, HRTC and Colorless showcase a $\Delta T_{30}$ lower than that of the proposed method, with an error of $22.6$ ms and $54.6$ ms against the $85$ ms obtained with our approach. Oppositely, both HRTC and the proposed FDN render the $T_{60}$ equally well, with errors of $10.1$ ms and $9.2$ ms, respectively. Our $\Delta C_{80}$ is one order of magnitude less than what can be achieved with the other methods. Likewise, our $\Delta D_{50}$ is one order of magnitude less than what can be achieved with GA and HRTC, and two orders of magnitude less than Colorless’. In addition, the center time error $\Delta t_{s}$ reaches $40.6$ µs, while all baselines have errors one to three orders of magnitude larger, thus shifting the center of mass of the predicted IR energy more toward the reverberation tail.
@@
Figure 12: Pizzeria (h214) IRs. The time axis is limited to the $T_{60}$ for visual clarity.
@@ 5.7 Test Case: Pizzeria (h214)
Finally, let us focus on the shortest RIR of the three considered in the present study, i.e., h214, having a $T_{60}$ of just above 0.2 s.
Here, all the baseline methods appear to fail at modeling the target room (Figs. 10 and 11). In particular, the IRs shown in Figure 12 exhibit a few sparse taps, being thus far from resembling a real RIR. Since the backward-integrated energy abruptly decays with every peak, this results in a staircase-like behavior in the early portion of the EDC depicted in Figure 10.
Facing the same difficulties, the proposed optimization method takes more gradient steps compared to the previous test case, converging at iteration 992 with a loss function two orders of magnitude lower than the starting value. In Figure 10, the resulting EDC (dashed orange line) closely follows that of the target RIR (solid black line) until the two reach approximately $-45$ dB. Still, a direct comparison between EDCs in the range of $-45$ to $-70$ dB is not entirely reliable since the RIR has so little energy that background and sensor noise take on a much more relevant role when it comes to integrating the energy of the measured signal.
Overall, the proposed optimization method proves to perform well in fitting the RIR under scrutiny, with Table 4 reporting an error of $4.7$ ms, $1.8$ ms, and $12.6$ ms when it comes to the three reverberation times. Furthermore, $\Delta C_{80}$ is 0.41 dB, $\Delta D_{50}$ is $0.13$%, and $\Delta t_{s}$ is $62.5$ µs. As in Section 5.5, the only metric for which the proposed method shows a performance worse than one of the baselines is the $T_{60}$. As a matter of fact, the HRTC is the only baseline to be once again characterized by a lower $\Delta T_{60}$, with a value of $5.2$ ms.
@@ 5.8 Excluding the EDP Loss Term
In developing our method, we noticed that using only the EDC loss function leads to ill-behaved IRs. Namely, we found that, while closely matching the desired EDC, the IR of an FDN trained with $\lambda=0$, i.e., using only $\mathcal{L}_{\text{EDC}}$, tends to exhibit an unrealistic echo distribution compared to the RIRs of real-life environments. In this section, we compare the results of our differentiable FDN trained without Soft EDP regularization with those presented in Section 5.6 obtained using the composite loss function in (8). For conciseness, we limit our analysis to the Hallway RIR (h270); results obtained with other RIRs are comparable to what is shown below.
When excluding the EDP loss term from (8), the NMSE between true and predicted EDCs (10) is comparable with that of the proposed method, totaling $2.8\times 10^{-3}$ ($\lambda=0$) and $3.4\times 10^{-3}$ ($\lambda=0.1$), respectively. Yet, the MSE between true and predicted EDPs (17) is $0.342$, i.e., two orders of magnitude higher than the $6.8\times 10^{-3}$ reported in Table 1 for the proposed method. This can be observed in Figure 13b, showing that the echo density when $\lambda=0$ (dash-dotted purple line) is far from the desired profile (solid black line). The target RIR and the proposed method (dashed orange line) produce an EDP with values consistently around one, indicating dense reverberation. On the contrary, the FDN trained without Soft EDP regularization yields an EDP with values below $0.5$, signaling an uneven echo density due to the presence of a few prominent reflections [39]. This observation is confirmed by the IR shown in Figure 13a that exhibits an exponentially decaying cluster of four taps periodically peaking above a denser reverberation tail with negative polarity.
Excluding the EDP loss term means solving a minimization problem with no constraints discouraging the model to use just a small number of delay lines to capture the overall EDC behavior. Since FDNs extend a parallel comb-filter structure, we believe that the behavior observed in Figure 13a is related to the well-understood problem occurring when only some delay lines are strongly excited and recirculated, which, in turn, aggravates the comb-like behavior of the delay network, ultimately resulting in an unpleasing metallic sound quality [42, 18].
Figure 13: (a) IR and (b) EDP obtained by training the FDN without EDP regularization term ($\lambda=0$).
@@ 5.9 Soft EDP Approximation
Finally, we discuss the approximation capabilities of the proposed Soft EDP function introduced in Section 4.5.2. Figure 14 shows the non-differentiable EDP (solid black line) defined in (11) against several Soft EDP approximations of the three RIRs considered in the present study. We test various scaling parameters $\kappa$, namely, $10^{2}$, $10^{3}$, and $10^{4}$, along with the proposed time-varying $\kappa_{n}$ linearly increasing from $10^{2}$ ($n=0$) to $10^{5}$ ($n=L_{T_{60}}-1$). We depict the profiles only for time indices below the $T_{60}$, as this range is the one considered when training the FDNs.
@@
HRTC $1.2566$ $0.3951$ $1.2597$ $0.3436$ $1.2194$ $\underline{0.0063}$
Colorless $1.2589$ $0.3973$ $1.2504$ $0.3342$ $1.1900$ $0.0357$
Ours $0.8451$ $\underline{0.0165}$ $0.9714$ $\underline{0.0552}$ $1.1355$ $0.0902$
@@
Target $12.106$ $89.009$ $22.899$
GA $11.985$ $0.1208$ $91.194$ $2.1848$ $19.343$ $3.5560$
HRTC $7.8183$ $4.2877$ $79.402$ $9.6070$ $38.346$ $15.447$
Colorless $2.7802$ $9.3258$ $56.329$ $32.679$ $77.035$ $54.135$
Ours $12.126$ $\underline{0.0200}$ $88.912$ $\underline{0.0974}$ $23.079$ $\underline{0.1805}$
Table 3: Metrics for the Hallway RIR.
@@
HRTC $0.6362$ $0.1072$ $0.6236$ $\underline{0.0226}$ $0.6168$ $0.0101$
Colorless $0.7007$ $0.1717$ $0.6556$ $0.0546$ $0.6335$ $0.0268$
Ours $0.4749$ $\underline{0.0540}$ $0.5160$ $0.0850$ $0.5975$ $\underline{0.0092}$
@@
Target $14.079$ $90.691$ $20.116$
GA $15.502$ $1.4223$ $92.106$ $1.4147$ $18.365$ $1.7514$
HRTC $11.412$ $2.6678$ $88.246$ $2.4453$ $20.793$ $0.6767$
Colorless $8.327$ $5.7520$ $77.850$ $12.841$ $31.269$ $11.152$
Ours $13.759$ $\underline{0.3204}$ $90.856$ $\underline{0.1648}$ $20.075$ $\underline{0.0406}$
Table 4: Metrics for the Pizzeria RIR.
@@
HRTC $0.2286$ $0.0643$ $0.2185$ $0.0391$ $0.2114$ $\underline{0.0052}$
Colorless $0.2668$ $0.1026$ $0.2557$ $0.0764$ $0.2367$ $0.0306$
Ours $0.1689$ $\underline{0.0047}$ $0.1811$ $\underline{0.0018}$ $0.1936$ $0.0126$
@@
Target $30.988$ $99.382$ $7.1698$
GA $40.683$ $9.6950$ $99.799$ $0.4170$ $6.4524$ $0.7174$
HRTC $24.860$ $6.1277$ $98.611$ $0.7708$ $9.4454$ $2.2756$
Colorless $21.849$ $9.1389$ $96.895$ $2.4862$ $10.749$ $3.5798$
Ours $30.576$ $\underline{0.4123}$ $99.250$ $\underline{0.1322}$ $7.2323$ $\underline{0.0625}$
@@ 5.10 Limitations
In this work, we focused on two important perceptual characteristics of room impulse responses: integrated energy decay and echo density. While rendering these time-domain features is key for any artificial reverberation algorithm that aims to be realistic, by themselves, Schroeder’s EDC and Abel and Huang’s EDP are not enough to comprehensively model the perceptual qualities of reverberation. In fact, it is well known that frequency- and time-frequency features play a crucial role in room acoustic simulation [4, 40].
However, Figures 17, 17, 17 show that none of the FDNs considered in the present study manages to capture the magnitude frequency response of the target RIRs. Likewise, Figure 18 reveals a significant discrepancy between the target Energy Decay Relief (EDR) [50] and those of the FDN models.
Such a conspicuous mismatch entails that the output signals of the FDNs sound different not only from one another but also from the corresponding input signal convolved with the target RIR. In turn, this undermines the reliability of any perceptual test assigning similarity scores to each method with respect to the target, as subjective judgments would be significantly influenced by differing spectro-temporal coloration and decay. In this respect, pilot experiments proved inconclusive, highlighting the need for further investigation into time-frequency modeling.
After all, matching the spectro-temporal characteristics of the target RIRs is not among the objectives of the parameter selection/tuning algorithms considered in the present study. Moreover, in this work, we mainly focused on time-invariant frequency-independent FDN prototypes. This holds true for proposed and baseline methods but GA, whose fitness function consists of a $L^{1}$-loss between MFCCs. Still, despite considering cepstral features, the result obtained with GA is far from resembling the reference spectro-temporal behavior. This evidences that it is not straightforward to accurately capture both time- and frequency-domain characteristics through an optimization process, even when including dedicated absorption and tone correction filters in the FDN prototype (cf. Figure 3).
Although said filters can be implemented in a differentiable fashion, the poor performance of GA points out that extending the cost function (8) to the frequency-dependent case may not be sufficient, ultimately suggesting that a more thorough and comprehensive study is necessary to accomplish the goal. We leave such an investigation for future work.
@@
Figure 18: Mel-frequency Energy Decay Reliefs (EDRs) of the Hallway test case (h270). Proposed in [50], the EDR extends Schroeder’s EDC to multiple frequency bands. Here, the 512-bin frequency axis is warped onto a 64-bin mel scale.